A survey conducted by USA Today showed that 118 of 250 investors owned some real estate.
Calculate a standard error

Answer by stanbon(57219)

A survey conducted by USA Today showed that 118 of 250 investors owned some real estate.
Calculate a standard error
----
p = 118/250 = 0.472 ; q = 0.528
-----
s = sqrt{p*q/n}
----
s = sqrt(0.472*0.528/250] = 0.0009969
================
Cheers,
Stan H.
================